# A Scene Representation for Online Spatial Sonification

### v1

@@ V Evaluation
We quantitatively evaluate the proposed representation for spatial sonification in terms of a) efficiency, b) representation accuracy, and c) comparison with Euclidean distance field, and d) qualitatively demonstrate the representation performance with a dynamic object. Our framework is implemented in C++ based on ROS1. All experiments were run on 12th Gen Intel® Core™ i5-1245U with 12 cores.
@@
The first dataset is the Cow and Lady[25], which includes fibreglass models of a large cow and a lady standing side by side in a room. The dataset consists of RGB-D point clouds collected using a Kinect 1 camera, with sensor trajectory captured using a Vicon motion system. We use all frames of the dataset, which covers major information of the scene over 360 degrees. We use the ground truth map of the Cow and Lady, which allows us to perform a proper quantitative evaluation for representation accuracy. We run the online mapping in a room with a live RealSense RGB-D camera as our second dataset. The room has a human as a dynamic object. The poses are computed from the VINS [26] localisation framework.
@@ V-A Time Evaluation
We first evaluate the time consumption to show our representation can provide online performance for sonification. Fig. 3 shows the statistical computational time for the Cow and Lady dataset. Note that our framework has the free-space carving methods [1] enabled to update the moving objects in the scene. For the sake of simplicity, we compute the time for all the processes, including fusion, 2D circle, and 3D cylinder. It illustrates that our representation is efficient with varying resolutions for online mapping.
@@ V-D Dynamic Moving Object Validation
We aim to show that our representation has the ability to deal with dynamic objects, which is very common in the real-world scenario. Our framework takes poses and points as raw input, so we use localisation frameworks such as VINs [26] to provide proper poses. We move a live camera to map the scene online. In Fig. 6, our 2D circle is fully reconstructed in 360 degrees, and we can visual object projection on the circle for each angle clearly. From Fig. LABEL:dynamic_1 to Fig. LABEL:dynamic_4, there was a human standing beside the sofa and later walked away. Our circle is updated with respect to the latest status of the scene.
Figure 7: a) Our circular representation shows the distance and directions along the ray. b)However, EDF shows the distance and directions to the closest surfaces.
@@ V-E Compare to EDF
We aim to show our representation is suitable for sonification rather than the Euclidean distance field (EDF), which is a well-used representation in path planning and navigation. For the EDF , it usually has a distance value and the gradient of the distance given the point’s location in the space. It always measures the distance and direction from the point to the closest surface. As we show in Fig.LABEL:arrow_ours, our circle has bearing angles pointing 360 degrees, and for each bearing, we have the closest distance along the ray (the ray arrow is shown in blue). However, as we see from Fig.LABEL:arrow_esdf, we query the EDF for each point on the circle, and this gives us the distance and gradient to the closest surfaces. The gradients of distances (shown in blue) on the circle slip around as the closest surface could be in any direction. This is not suitable for sonification as the bearing does not vary sequentially.
@@ V-F Sonification Demonstration
In our accompanying video [^1] https://tinyurl.com/RepresentationSonification, we show the performance of our proposed representation for sonification with both modes, circular ranging and circular ranging of objects, in the room-size environments of our datasets.
@@ VI Conclusion
We propose an online, incrementally built representation to provide spatial auditory information to aid individuals with visual impairments. By leveraging a sensor-centric mapping structure based on depth sensors, we effectively convert 3D environments into 360-degree auditory representations. Our approach, which includes the VDB-GPDF, plane projection, rasterisation and binaural room impulse responses, ensures fast accurate, complete and perceptually robust mapping for sonification. We demonstrate its performance in both static and dynamic environments. The successful implementation and evaluation of this framework highlight its potential to significantly improve the quality of life for visually impaired individuals by providing an intuitive and efficient means of navigating their surroundings. In future work we aim to perform user studies targeting to reduce the cognitive load of the users.

### v2

@@ V Evaluation
To the best of our knowledge, no open-source state-of-the-art spatial sonification frameworks are available for comparison. Therefore, we quantitatively evaluate the proposed representation for spatial sonification in terms of a) efficiency, b) representation accuracy and c) coverage with respect to using depth images, d) qualitatively demonstrate the representation performance with a dynamic object and e) suitability for sonification with respect to Euclidean distance fields. Our framework is implemented in C++ based on ROS1. All experiments were run on 12th Gen Intel® Core™ i5-1245U with 12 cores.
The first dataset is the Cow and Lady[25], which includes fibreglass models of a large cow and a lady standing side by side in a room. The dataset consists of RGB-D point clouds collected using a Kinect 1 camera, with sensor trajectory captured using a Vicon motion system. We use all frames of the dataset, which covers major information of the scene over 360 degrees. We use the ground truth map of the Cow and Lady, which allows us to perform a proper quantitative evaluation for representation accuracy. As our second dataset, we run the online SLAM in a room with a live RealSense RGB-D camera. The room has a human as a dynamic object. The poses are computed from the VINS [4] localisation framework as described above.
@@ V-A Time Evaluation
We first evaluate the time consumption to show our representation can provide online performance for sonification. Fig. 4 shows the statistical computational time for the Cow and Lady dataset. Note that our framework has the free-space carving methods [1] enabled to update the moving objects in the scene. For the sake of simplicity, we compute the time for all the processes, including fusion, 2D circle, and 3D cylinder. It illustrates that our representation is efficient with varying resolutions for online performance.
@@ V-D Dynamic Moving Object Validation
We aim to show that our representation has the ability to deal with dynamic objects, which is very common in the real-world scenario. Our framework takes poses and points as raw input, so we use localisation frameworks such as VINS [4] to provide proper poses. We move a live camera to map the scene online. In Fig. 7, our 2D circle is fully reconstructed in 360 degrees, and we can visual object projection on the circle for each angle clearly. From Fig. LABEL:dynamic_1 to Fig. LABEL:dynamic_4, there was an object in the scene and later moved away. Our circle is updated with respect to the latest status of the scene.
Figure 8: a) Our circular representation encodes the radial distance at a given angle over the scan. b) EDF, however, encodes the distance and directions to the closest surfaces.
@@ V-E Compare to EDF
This section aims to show the suitability of our representation for sonification over the Euclidean distance fields (EDF), which is better suited for path planning and navigation. For the EDF [1], it usually has a distance value and the gradient of the distance given the point’s location in the space. It always measures the distance and direction from the point to the closest surface. As we show in Fig. LABEL:arrow_ours, our circle has bearing angles pointing 360 degrees, and for each bearing, we have the closest distance along the ray (the ray arrow is shown in blue). However, as we see from Fig. LABEL:arrow_esdf, we query the EDF for each point on the circle, and this gives us the distance and gradient to the closest surfaces. The gradients of distances (shown in blue) on the circle slip around as the closest surface could be in any direction. This is not suitable for sonification as the bearing does not vary sequentially.
Figure 9: Incrementally built scene with a live depth camera following our framework mapping, self-localising, and maintaining the proposed sonification representation.
@@ V-F Sonification Demonstration
In our accompanying video and Fig. 9, we show an incrementally built and sonified room-sized environment using the proposed framework. Incremental localisation (using VINS [4]), mapping (with VDB-GPDF [1]), circular rasterization and sonification are shown in real-time. Listening through headphones is recommended to hear the sounds at the correct azimuth location.
@@ VI Conclusion
We propose an online, incrementally built representation to provide spatial auditory information of unknown environments. This approach has the potential to aid individuals with visual impairments. By leveraging a sensor-centric mapping structure based on depth sensors, we effectively convert 3D environments into 360-degree auditory representations. Our approach, which includes the VDB-GPDF, plane projection, rasterisation and binaural room impulse responses, ensures fast accurate, complete and perceptually robust mapping for sonification. We demonstrate its performance in both static and dynamic environments. The successful implementation and evaluation of this framework highlight its potential to significantly improve the quality of life for visually impaired individuals by providing an intuitive and efficient means of navigating their surroundings. In future work, we aim to perform user studies targeting to reduce the cognitive load of the users for the 3D representation.